OpenStudy (anonymous):

Find an equation of the tangent line to the curve y=2x sin x at the point (pi/2,pi)

OpenStudy (raden):

first, find the slope (m) = y'

OpenStudy (anonymous):

find dy/dx at x=pi/2

OpenStudy (anonymous):

Wouldn't that be 2cos x? and if you plug in pi/2 or 90 degrees you get 0?

OpenStudy (raden):

use the product rule y ' = vu' + uv"

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.

OpenStudy (anonymous):

then eq. of tangent is \[y-\pi=(\frac{ dy }{dx }at x=\frac{ \pi }{2 })\left( x-\frac{ \pi }{2 } \right)\]

OpenStudy (anonymous):

Okay after I do the product rule, which I end up with 2, what step do I take from there?

OpenStudy (anonymous):

And after plugging in x*

OpenStudy (raden):

yes, correct m = 2 then use the equation like @surjithayer said above :

OpenStudy (anonymous):

Phew thank you! While Im here can you I ask of you for help on another problem?

OpenStudy (raden):

what's the question then ..

OpenStudy (anonymous):

Differentiate. f(x)= xe^x csc x

OpenStudy (raden):

let u = xe^x and v = cscx use the product rule

OpenStudy (raden):

to get u' looks we need the product rule again :)

OpenStudy (raden):

u = xe^x u ' = (xe^x)' = x' e^x + x (e^x)' = 1 e^x + x e^x = e^x + xe^x, agree ?

OpenStudy (anonymous):

Oooh okay, haha yeah I was just confusing myself I agree yes, Thank you for your help!

OpenStudy (raden):

oke, go back the original question :) f(x)= xe^x csc x f '(x) = (xe^x)'csc x + xe^x(csc x)' (xe^x)' already done above, now (cscx)' = -cscxcotx so, f ' = u'v + uv' = ....

OpenStudy (anonymous):

xe^xcsc + e^x cscx-xe^x cscx-xe^xcscx + xe^x cotx, and I can see that you can just factor out an xe^xcsc so final answer would be.. e^xcscx(-x cot x +x +1) ?